# Benchmarking Performance, Explainability, and Evaluation Strategies of Vision-Language Models for Surgery: Challenges and Opportunities

### v1

@@ Abstract
Minimally invasive surgery (MIS) presents significant visual and technical challenges, including surgical instrument classification, and understanding surgical action involving instruments, verbs and anatomical targets. While many machine learning-based methods have been developed for surgical understanding, they typically rely on procedure and task-specific models trained on small, manually annotated datasets. In contrast, the recent success of vision-language models (VLMs)—models trained on large volumes of raw image-text pairs—has demonstrated strong adaptability to diverse visual data and a range of downstream tasks. This opens meaningful research questions: how well do these general-purpose VLMs perform in the surgical domain, In this work, we aim to explore those questions by benchmarking several VLMs across diverse surgical datasets, including general laparoscopic procedure and endoscopic submucosal dissection to assess their current capabilities and limitations. Our benchmark work reveals key gaps in the models’ ability to correctly link language to the right areas in the surgical scene in a consistent way.
@@ 1 Introduction
Minimally invasive surgery (MIS) offers substantial benefits for patients, including smaller incisions, reduced bleeding and pain, and faster recovery times. However, MIS remains technically challenging for surgeons due to a limited field of view, reduced depth perception, and constrained hand-eye coordination. To address these challenges, many studies have proposed machine learning algorithms for classifying and segmenting surgical instruments and tissues, recognizing procedural phases, and assessing surgical skill Twinanda et al. (2016a); Jin et al. (2020); Zisimopoulos et al. (2018). This approach helps interpret surgical progress, inform potential risks and support real time assistant for surgeons. Despite these efforts, most existing approaches still rely on training task-specific models using relatively small, manually labeled datasets created by clinical experts. These datasets are costly to annotate and limited in scale, which hinders generalization. Moreover, surgical scenes present inherent visual challenges, including varying instrument distances and poses, deformable and visually similar tissues, occlusions, motion blur, and poor lighting conditions.
Minimally invasive surgery (MIS) is routinely performed under video guidance, meaning surgeons rely heavily on visual feeds to carry out procedures. As a result, large volumes of surgical video data are generated daily, much of which already exists but remains underutilized for computational analysis and model training. Additionally, numerous instructional and teaching materials are publicly available online. Identifying the value of raw surgical data remains an open question, particularly in how to effectively leverage them. Recently, the success of vision-language models (VLMs) has demonstrated the potential of learning from large-scale, unstructured data. Models like CLIP Radford et al. (2021) learn visual perception from supervision embedded in natural language, using paired images and text instead of meticulously labeled datasets. This paradigm aligns naturally with the use of raw surgical data and opens promising avenues for surgical AI. When paired with audio, reports or lecture material, large-scale surgical video recording could help overcome the limitations of current models that are restricted to small, manually annotated datasets. Moreover, the revolution in natural language processing enables a more comprehensive understanding of the surgical environment, far beyond simple classification tasks. VLMs have demonstrated remarkable capabilities in processing and reasoning over both visual (e.g., images, videos) and textual (language) data. By combining computer vision and natural language, they allow machines to interpret, reason, and align across modalities Radford et al. (2021); Li et al. (2022); Yuan et al. (2024a, b, 2023).
Figure 1: Examples of VLM doing surgical instrument classification task using CLIP-ViT-B/32 with Grad-CAM to visualize where VLM is looking at. Images (a) and (b) are sampled from the CholecT80-BoundingBox dataset Twinanda et al. (2016b), while (c) and (d) are from the CoPESD dataset Wang et al. (2024). Image pairs (a) and (b) illustrate common spatial reasoning challenges in surgical environments: instruments may appear at varying distances from the endoscope, partially occluded, or positioned at non-standard angles. Image pairs (c) and (d) highlight the effects of poor lighting conditions and motion blurring, where limited illumination leads the model to focus on irrelevant or background regions.
In scenarios involving surgical robots or AI assistants, it is crucial to detect when a model may fail so that control can be safely handed back to the surgeon. Explainability is especially important. Knowing when and how the model fails matters, as the consequences of failure are significantly higher. This requires the model not only to process complex scenes, but also to focus on the most relevant visual cues. Recent work by Wang et al. Wang et al. (2023) highlights a key concern in vision-language models (VLMs): they often focus on irrelevant background regions rather than task-relevant objects, raising concerns about models making correct predictions for the wrong reasons and undermines model’s reliability and interoperability. As shown in Figure 1.
Given that the application of VLMs in surgical environments is still relatively underexplored. With variety of available VLM architecture and strategy for training, and substantial resources required to collect, this leads to a key question that we aim to explore through benchmarking: How effectively do current vision-language models perform on surgical data? To the best of our knowledge only one work by Rau et al. Rau et al. (2025) has benchmarked VLMs across tasks like classification, detection, segmentation, and risk assessment, these evaluations primarily rely on conventional metrics such as accuracy, F1 score, and mIoU. These metrics fail to capture whether the model’s predictions are grounded in clinically meaningful reasoning. To address this gap, we propose a framework centered on examining model attention during prediction, revealing not only what the model predicts but also whether it attends to the correct regions in the image. Specifically, we incorporate class activation mapping (CAM) techniques such as Grad-CAM Selvaraju et al. (2017), which highlight the spatial regions that contribute most to a model’s decision. These heatmaps allow us to directly assess whether the model’s attention is aligned with the relevant surgical content. By visualizing and quantifying the alignment between model attention and ground-truth regions, in this benchmark work, our main contributions are as follows: (1) We introduce a benchmarking framework for evaluating vision-language models (VLMs) in surgical settings, with a focus on the attention of VLMs beyond conventional classification accuracy. (2) We propose multiple new metrics that quantify whether the focus of a model aligns with relevant regions in the surgical data. (3) We demonstrate a consistent gap between correct predictions and correct attention, raising concerns about the reliability of VLMs in safety-critical surgical contexts.
@@ 2.1 VLM
We benchmark pre-trained vision-language models (VLMs) on tasks involving surgical scene understanding, with a focus on their zero-shot interpretability and spatial alignment. Prior work has employed zero-shot approaches to leverage pretrained models for downstream tasks without the need for fine-tuning, thereby significantly reducing computational resource requirements Alayrac et al. (2022). A notable example is CLIP Radford et al. (2021), which learns aligned image-text representations from large-scale web data and has been widely used in zero-shot classification and retrieval tasks. Techniques such as prompt learning and tuning strategies have also been proposed to improve visual-textual embedding alignment in VLMs Liu et al. (2023); Zhou et al. (2022); Chen et al. (2025). In this work, we adopt a benchmarking approach using frozen models, specifically evaluating their zero-shot generalization capability in the surgical domain.
@@ 2.2 VLM for surgical applications
There is growing attention toward joint vision-language learning in surgical domains. In reality, only a handful of studies have begun to explore this area learning joint representations from surgical videos and natural language descriptions, often sourced from both narrated and slide-guided surgical lectures, to obtain more generalization VLMs. SurgVLP Yuan et al. (2023) introduced a large-scale surgical pretraining approach using data from surgical lecture videos. It adopts a CLIP-style training paradigm with a contrastive learning objective to align video clips and their corresponding textual descriptions in a shared embedding space. A follow-up model, HecVL Yuan et al. (2024b), enhances this framework by introducing hierarchical supervision, pairing videos with both free-form text and structured phase- and procedure-level annotations, leading to improved phase recognition. PeskaVLP Yuan et al. (2024a) further extends this line of work by enabling alignment between video and text even when they are not temporally synchronized. Other foundation model such as GSViT Schmidgall et al. (2024), which is also trained on surgical video data using a video prediction objective. While GSViT effectively captures temporal dynamics and achieves strong performance in surgical phase recognition. Most prior work applying VLMs to surgical tasks has focused on temporal modeling and phase recognition. Through our benchmark, we aim to systematically evaluate these models from a different perspective—specifically, their spatial and semantic grounding abilities in zero-shot settings. Rather than replacing prior approaches, our work complements them by identifying limitations across spatial, semantic dimensions. This helps inform future efforts to better adapt and align VLMs with the unique demands of surgical applications.
@@ 2.3 Visual explanation methods for Deep learning
While deep nerual networks enable superior performance their lack of decomposability into intuitive and understandable components makes them hard to interpret. Consequently, when today’s intelligent systems fail, they fail spectacularly, disgracefully, without warning or explanation, leaving a user staring at an incoherent output. Several methods have been developed to visualize which regions of an image have the most influence on model predictions. Grad-CAM Selvaraju et al. (2017) highlights important regions by computing the gradients of the output with respect to intermediate feature maps of convolutional networks. For transformer-based models, recent work Chefer et al. (2021) extends this idea by computing gradients with respect to attention layers. Grad-CAM++ Chattopadhay et al. (2018) refines the original approach by improving localization in cases with multiple object instances, using a weighted average of higher-order derivatives. Score-CAM Wang et al. (2020), on the other hand, avoids gradient computation altogether by using the model’s output scores to directly weight the activation maps. This can lead to more stable results, though at a higher computational cost. In our work, we adopt the original Grad-CAM method due to its simplicity and ease of implementation, and use it as a core tool in our benchmark.
@@ 2.4 Evaluation metrics for Surgical data
Recent benchmark papers on VLM understanding in surgical environments have focused on the general performance of VLMs. Rau et al. (2025) primarily evaluating classification results for tasks such as tool and anatomy recognition, segmentation, and identifying risk or safety-related information. However, conventional metrics like accuracy or IoU often fail to capture whether the model’s predictions are based on clinically meaningful evidence, especially in high-stakes surgical contexts. To our knowledge, the only work that explicitly questions the suitability of conventional metrics in surgical data is Funke et al. (2023), which focuses on the task of surgical phase recognition and highlights that conventional metrics fail to capture temporal structure such as the correct order and duration of phases. For example, a model that predicts all the right phases but in the wrong sequence, or switches phases too frequently. To address this, they propose using metrics like segmental F1 score and edit distance, which evaluate how well the predicted phase segments match the ground truth in terms of both timing and order. These metrics penalize mistakes such as switching phases too often or predicting the correct phases in the wrong sequence, which are better suited to evaluate whether the model predicts the correct sequence and timing of surgical phases. For quantifying spatial semantic grounding, Chen et al. (2022) use the idea of IoU to evaluate gScoreCAM by measuring the spatial overlap between the model’s binarized attention region and the ground-truth object. However, this approach overlooks how strongly different regions contribute to the prediction, which may carry important information in surgical data due to the complexity of the visual scene. Another limitation is that if the model’s prediction is incorrect, its attention is likely to be misaligned as well—making it difficult to judge whether the attention mechanism is meaningful or reliable when compared to ground-truth regions.
@@ 3 Proposed Methods
In this section, we first provide a brief introduction to the goal of this benchmark work in Section 3.1. We then describe conventional tasks on tasks this benchmark will evaluate and Class Activation Mapping (CAM) techniques used to interpret VLMs, including methods designed for ResNet-based architectures as well as those adapted for transformer-based models. Finally, we describe the design of our metrics, which aim to quantify the spatial reasoning capabilities of VLMs in surgical environments and provide straightforward visualizations.
3.1 Problem and Task Definition
In this paper, we examine whether VLM predictions in surgical settings are grounded in the correct visual evidence. Formally, this motivates the following question: Given an image $I\in\mathbb{R}^{H\times W\times 3}$ and a set of natural language prompts $\mathcal{T}={T_{1},T_{2},\dots,T_{n}}$, does the model’s attention align with the image regions in $I$ that are relevant to the predicted prompt? Answering this requires going beyond classification accuracy and assessing whether the model’s visual focus meaningfully supports its predictions. To do so, we first introduce visualization-based methods that evaluate this alignment. Our study focuses on three key tasks: (i) instrument classification and grounding and (ii) surgical action classification via triplet descriptions. These tasks evaluate different aspects of multimodal model understanding, including spatial grounding, contextual reasoning, and object/background discrimination in surgical environments.
@@ 3.1 Problem and Task Definition
In the instrument classification task, a vision-language model (VLM) selects the most semantically aligned description from a pool of prompts, each referring to a specific surgical instrument class. Conventional evaluation for this task simply measures the proportion of frames where the predicted instrument matches the ground truth label.
@@
In Triplet classification task, vision-language model (VLM) selects the most semantically aligned action description from a pool of candidate triplet prompts to evaluate the model’s ability to infer structured surgical actions. Each prompt follows the template “I use a [instrument] to [verb] the [target]”, representing subject-action-object combinations.Conventional evaluation metrics for this task are similar to the instrument classification task and straightforward, which measures how often each individual component is predicted correctly. Let $x\in\{s^{*},v^{*},o^{*}\}$, then the accuracy for component $x$ is computed as the proportion of frames where the predicted $x$ matches the ground truth.
3.2 Proposed CAM-based metrics
Class Activation Mapping (CAM) methods aim to localize the regions in an image that contribute to a model’s decision. These methods work by analyzing the internal activations of a convolutional neural network and their influence on a target output score. We first talk about how Grad-Cam work in VLM that use ResNet.

### v2

@@ Abstract
Minimally invasive surgery presents significant visual challenges, including a limited field of view, specular reflections, and inconsistent lighting conditions due to the small incision and use of endoscopes. Over the past decade, while many machine learning and deep learning models have been developed to identify and detect instruments and anatomical structures in surgical videos, these models are typically trained on manually labeled, procedure- and task-specific datasets that are relatively small, results in limited generalization capability to unseen data. In fact, a massive amount of raw surgical data, including videos captured during various surgical procedures everyday in hospitals. Given that labeling all of this raw surgical data is nearly impractical due to the need for highly specialized expertise, recent success of vision-language models (VLMs) - which can be trained on large volumes of raw image-text pairs and gain strong adaptability - offers a promising alternative for leveraging raw surgical data. While a few studies have begun exploring VLMs for surgical data, their performance remains limited. To support future development of more effective VLMs for surgical applications, this paper aims to answer a key question: How well do existing VLMs - both general-purpose and surgery-specific - perform on surgical data, and what type of scenes do they struggle with? Therefore, we conduct a benchmarking study of several popular VLMs across diverse datasets of gastrointestinal surgery and general surgery. Specifically, we visualize the model’s attention to identify which regions of the image it focuses on when making predictions for surgical tasks. We also propose several metrics to evaluate whether the model attends to task-relevant regions. Our findings reveal a mismatch between prediction accuracy and visual grounding.
@@ 1 Introduction
Minimally invasive surgery (MIS) enables procedures to be performed through small incisions, offering substantial benefits for patients such as reduced bleeding and pain, and faster recovery times. However, MIS also brings many challenges for surgeons including the limited field-of-view and reduced depth perception from the endoscope, which is a small camera inserted into the patient body to provide visualization of the surgical site, affecting surgeons’ hand-eye coordination ability [1]. To reduce unintended damage and enhance patient outcomes, many studies have proposed machine learning (ML) and deep learning (DL) models to analyze surgical videos, with the goal of providing real-time guidance and informing potential risks to surgeons during procedures. These models have been applied to tasks such as surgical tool and tissue classification, semantic and instance segmentation, phase and action recognition, workflow analysis, and surgical skill assessment [2, 3, 4]. Despite these efforts, most existing models in this field remain task-specific and were trained using relatively small datasets, usually consisting of thousands of images manually annotated by clinical experts. However, due to visual challenges inherent in surgical scenes, such as varying lighting conditions, textureless and moist tissues, occlusions, motion blur, and the presence of blood and smoke, these datasets are often insufficient for models to generalize effectively to unseen videos, even within the same type of procedure.
Although manual annotation of surgical data is difficult to scale, if a model can tolerate low-quality labels, there actually exists a massive amount of raw surgical data that can be leveraged. For example, since surgeons cannot view the surgical site directly as in traditional open surgeries, MIS has been routinely performed under video guidance. As a result, large volumes of surgical video data are generated every day, much of which is already stored in hospital databases. Additionally, numerous instructional videos and educational resources for medical education are publicly available online. If we can develop a model capable of efficiently learning from such large-scale raw surgical data, it may achieve a level of generalization that is hard to reach using existing ML/DL models for surgical applications trained on much smaller datasets. Models like CLIP [5] learn visual-text representations using contrastive learning, a technique that encourages the model to bring representations of paired images and text closer together in a multi-modal embedding space, while pushing apart representations of mismatched pairs. This approach reduced the need for meticulously labeled datasets by leveraging weak supervision from natural correspondences between images and their descriptions. The success of VLMs points out a promising way to use large volumes of raw surgical data.
Figure 1: Examples of a VLM, CLIP-ViT-B/32, performing surgical instrument classification. The first column shows two input images. The next two columns show the classification results from the VLM and the corresponding model attention maps generated by Grad-CAM. We report the cosine similarity scores estimated by the VLM for the top two classes with highest cosine similarity scores. Green bounding boxes overlaid on the attention maps are ground truth annotations, (a) from the CoPESD dataset [11] and (b) from the CholecT80-BBox dataset [10] both show that VLMs may not consistently focus on the most relevant locations.
Recent efforts in surgical AI have begun to explore the development of foundation models for surgical environments [35]. SurgVLP, HecVL, and Peskavlp[9, 8, 7] are trained using contrastive learning on surgical videos paired with text, with each method adding progressively more structure or supervision to improve alignment. Moreover, surgical data exhibits a long-tailed distribution, with many rare yet complex cases appearing only occasionally during procedures [29, 31]. As a result, even models with strong generalization ability, including such as VLMs, can still fail in these less-represented scenarios.
In surgery, a single misjudgment can lead to serious complications like excessive bleeding, organ damage, or even life-threatening outcomes. Therefore, it is crucial to detect when a model may fail so that clinicians are aware of the limitations of this model and know when to reply more on their own expertise to make a decision. Explainability in AI is an evolving field focused on interpreting model predictions and understanding when and how the model fails. Explainability requires the model not only to make accurate prediction but also to ground it reasoning the most relevant cues. Given the high-stakes nature of surgical applications, incorporating explainability into these models is essential for ensuring informed clinical decision-making. Recently, there has been growing interest in providing explanations for neural networks to better understand their transparency and credibility [12, 21, 24, 20, 30]. In particular, recent work [12] highlights a key concern in VLMs: they often attend to irrelevant background regions rather than task-relevant objects, raising concerns about models making correct predictions for the wrong reasons, which undermines their reliability and interoperability. In this work, we adopt existing explainability techniques to study the reliability and interpretability of VLMs in surgical settings. Specifically, we incorporate CAM methods such as Grad-CAM [14] to visualize the spatial regions that contribute most to a model’s prediction. These heatmaps allow us to assess whether the model’s attention aligns with the relevant surgical content, as illustrated in Figure 1.
Given that the application of VLMs in surgical environments is still in its early stages, and considering the diversity of available VLM architectures, training strategies, and the substantial resources required for data collection, curation, model training, a benchmarking study is needed to evaluate the performance and limitations of existing VLMs and to guide future development of surgical VLMs. To the best of our knowledge, only one prior work by Rau et al. [13] has conducted a comprehensive benchmark of VLMs in the surgical domain, evaluating 11 state-of-the-art VLMs across 17 core surgical visual understanding tasks, including classification, detection, segmentation, and risk assessment. Their results indicate that while current VLMs demonstrate strong generalization abilities, their performance remains far below the accuracy required for practical application in surgical tasks. However, their evaluation focuses primarily on standard metrics such as F1 score, revealing only the average performance across datasets, which may not provide sufficient insight for determining effective surgical VLM training strategies. Therefore, in this work, we adopt an existing neural network explainability technique CAM to analyze which image regions have greater impact on the model’s decision-making process, thereby providing better guidance for future development of surgical VLMs. In this work, we benchmark pre-trained VLMs on tasks involving surgical scene understanding, with a focus on their zero-shot interpretability and spatial alignment. Our main contributions are as follows:
  • We introduce a VLM benchmarking framework for surgical data that goes beyond standard evaluation workflows by incorporating model explainability. We visualize attention maps to reveal which regions of an image have the greatest influence on the model’s predictions.
  • We propose several new metrics to quantitatively analyze the attention maps and evaluate the reliability of VLM predictions. These metrics complement standard metrics and enable a more comprehensive evaluation of the models
@@ 2.1 VLM
VLMs have established a powerful paradigm for bridging visual and textual modalities. A prominent example is CLIP [5], which learns visual-text representations through contrastive learning using an image encoder (e.g., ResNet [33] or Vision Transformer [32]) and a text encoder that learn a joint embedding space by aligning images and natural language descriptions during training on large-scale image-text pairs from the web. This approach leverages available web data without requiring manual annotations, demonstrating how models can learn meaningful representations directly from raw, unstructured data sources. At inference time, CLIP takes an image and a set of natural language prompts (e.g., “a photo of a grasper” or “an image showing surgical scissors”) and computes the similarity between the image embedding and each text embedding to determine the most relevant match. The effectiveness of CLIP-style VLMs heavily depend on how well the text prompts align semantically and structurally with the image content. which ideally should imply that the VLM can focus on the correct regions of the image that correspond to the described concepts.
@@ 2.2 VLM for surgical applications
Recent efforts on VLMs for surgical applications have explored using training data from both narrated surgical videos and slide-guided surgical lectures to enhance the models’ generalization capabilities. SurgVLP [9] introduced a large-scale surgical pretraining approach using data from surgical lecture videos. It adopts a CLIP-style training paradigm with a contrastive learning objective to align video clips and textual descriptions in a shared embedding space. A follow-up model, HecVL [8], enhances this framework by introducing hierarchical supervision, pairing videos with both free-form text and structured phase- and procedure-level annotations, leading to improved phase recognition. PeskaVLP [7] further extends this line of work by incorporating LLM-augmented surgical knowledge into the texts and applying a procedure-aware loss to improve cross-modal temporal alignment . Other foundation model such as GSViT[36], trained on 680 hours of surgical videos that learns to predict what happens next in surgery videos, and is fast enough to analyze surgical footage in real-time to potentially assist surgeons or control surgical robots during live operations. Learning from a large number of surgical videos and lectures allows surgical VLMs to have the ability to recognize surgical phases and identify the actions performed in different video clips. In this work, we also evaluate these models’ performance on more detailed object-level understanding, such as recognizing tissues and instruments.
@@ 2.3 Visual explanation methods for Deep learning
While deep learning enable superior performance, their lack of decomposability into intuitive and understandable components makes them hard to interpret. Consequently, when today’s intelligent systems fail, they fail without warning or explanation, producing outputs that cannot always be trusted. To mitigate this challenge, several methods have been developed to visualize which regions of an image have the most influence on model predictions, with the goal of revealing the evidence the model relies on to make its predictions, thereby providing insight into their explainability. Grad-CAM [14] highlights important regions by computing the gradients of the output with respect to intermediate feature maps of convolutional networks, and demonstrates its usefulness in helping users build appropriate trust in model predictions. Recent work [20] extends this idea by computing gradients with respect to attention layers. Other works such as, Grad-CAM++ [21] refines the original approach by improving localization in cases with multiple object instances, using a weighted average of higher-order derivatives. Score-CAM [22], on the other hand, avoids gradient computation altogether by using the model’s output scores to directly weight the activation maps. This can lead to more stable results, though at a higher computational cost. In our work, we adapt the original Grad-CAM method due to its simplicity and ease of implementation, and use it as one core tool in our benchmark.
@@ 2.4 Evaluation metrics for Surgical data
To the best of our knowledge, Rau et al. [13] is the only prior work that conduct benchmarking of VLMs on surgical data. It evaluates several commonly studied tasks such as instrument and anatomy classification, segmentation, and the identification of risk- or safety-related information, using standard evaluation metrics like the F1 score. Recently, there are studies highlights the need for new metrics to complement standard evaluation methods by capturing aspects of model performance that traditional metrics may overlook [23]. Given that our paper focuses on analyzing model attention, we propose several new metrics to assess the extent to which the model’s predictions rely on correct attention. These metrics are inspired by several prior studies that analyze attention maps to gain insight into model behavior. For spatial semantic grounding, Chen et al. [24] introduce the use of IoU to evaluate gScoreCAM by measuring the overlap between binarized attention maps and ground-truth object regions. However, this approach overlooks the varying strength of contribution across different regions—an important factor in surgical images due to their complex visual scenes. Furthermore, when the model’s prediction is incorrect, its attention is often misaligned, making it difficult to assess whether the attention mechanism is meaningful or trustworthy. Choe et al. [37] propose the Localization Accuracy metric for weakly supervised object localization, which considers a prediction correct only when both the predicted class label and the localized region (from a saliency map) align with the ground truth. It evaluate not only attending to the correct region but also predicting the correct semantic label.
@@ 3 Proposed Methods
In this section, we first introduce the visual tasks for understanding surgical images in Section 3.1 and problem evaluated in this benchmark in Section 3.2. We then describe the Class Activation Mapping (CAM) techniques, including methods designed for ResNet-based architectures as well as those adapted for transformer-based models in Section 3.3. Finally, we describe the metrics used to evaluate the VLMs, including standard task metrics as well as newly proposed metrics for analyzing attention maps, as described in section 3.4.
3.1 Task Definition
Our study focuses on two key tasks: (i) instrument classification and grounding and (ii) surgical action classification via triplet descriptions.
@@ 3.1 Task Definition
In the instrument classification task, a VLM selects the most semantically aligned description from a pool of prompts, each referring to a specific surgical instrument class.
@@
In the triplet classification task [34], a VLM is evaluated on its ability to infer structured surgical actions represented as subject-action-object triplets,where the subject denotes the surgical instrument (e.g., grasper), the action refers to the surgical motion (e.g., retract), and the object indicates the anatomical target (e.g., gallbladder).
3.2 Problem Definition
In addition to benchmarking the performance of VLMs on surgical data, a main objective of this work is to assess whether their predictions are grounded in the correct visual evidence. Formally, this motivates the following question:
  • When the model makes a correct prediction, does its attention focus on the relevant regions in the image that support that decision?
  • When the model makes an incorrect prediction, is it attending to the wrong visual cues?
By examining these two scenarios, we aim to better understand not just what the model predicts, but why—providing insight into its reasoning process and the reliability of its visual grounding. Answering this requires going beyond classification accuracy and assessing whether the model’s visual focus meaningfully supports its predictions. To do so, we introduce visualization-based methods in the next section.
3.3 Class Activation Mapping
Class Activation Mapping (CAM) methods aim to localize the regions in an image that contribute to a model’s decision. These methods work by analyzing the internal activations of a convolutional neural network and their influence on a target output score. Here, we provide a brief explanation of how Grad-CAM works in two commonly used visual encoders of VLMs: ResNet[33]and Visual Transformer[32].